A postal clerk sold 70 stamps for $17.00.Some were 35 cent stamps and some were 20 cent.How many of each stamps did he sell?
Answer by checkley71(8403)

X+Y=70 & 35X+20Y=1700 THUS X=70-Y SUBSTITUTING (70-Y) FOR X IN THE SECOND EQUATION WE GET 35(70-Y)=20Y=1700 OR 2450-35Y+20Y=1700 OR -15Y=-750 OR
Y=750/15 OR Y=50 THEREFORE X=70-50 OR X=20
PROOF 35*20+20*50=1700 OR 700+1000=1700